OpenStudy (anonymous):

Write a balanced equation showing both the first and second dissociation reactions for carbonic acid, H2CO3.

sam (.sam.):

Carbonic acid is diprotic: it has two protons, which may dissociate from the parent molecule. Thus, there are two dissociation constants, the first one for the dissociation into the bicarbonate (also called hydrogen carbonate) ion HCO3−: \[H_2CO_3\leftrightharpoons HCO_3^− + H^+\] \[K_{a1} = 4.5×10−7 ~~;~~ pK^{a1} = 6.367~at~25 °C\] The second for the dissociation of the bicarbonate ion into the carbonate ion \[CO_3^{2−}:\] \[HCO^{3−} \leftrightharpoons CO_3^{2−} + H^+ \] \[K_{a2} = 4.69×10−11 ~;~ pK_{a2} = 10.329 ~at~ 25 °C ~and~ Ionic~ Strength~ =~ 0.0 \]